OpenStudy (anonymous):

2(3)^2x=5 solve. on a page about logarithms

jimthompson5910 (jim_thompson5910):

\[\large 2(3)^{2x}=5\] \[\large (3)^{2x}=\frac{5}{2}\] \[\large 2x=\log_{3}\left(\frac{5}{2}\right)\] \[\large x=\frac{\log_{3}\left(\frac{5}{2}\right)}{2}\]